Solve the equation 25^(x+1)=125^x
Help!

Answer by venugopalramana(3286)

LET 5^X=Y
> 25^X=(5^2)^X=5^2X=(5^X)^2=Y^2
> 125^X=(5^3)^X=5^3X=(5^X)^3=Y^3
WE HAVE....25^(X+1)=125^X...
(25^X)(25)=125^X...OR...
25Y^2=Y^3
Y^2(Y-25)=0...OR...Y=25..
5^X=25...SO..X=2